OpenStudy (cutiecomittee123):

Solve for x 150=50e^7x

OpenStudy (cutiecomittee123):

a. x=ln(3)-7 b. x=l/e c. x=ln(3)/7 d. x=ln(150)/79(ln)(50)

OpenStudy (cutiecomittee123):

I think its c, right?

OpenStudy (campbell_st):

same process as before divide both side by 50 then take the logs and use the rule for powers

OpenStudy (cutiecomittee123):

Okay I divided by 50 and got 3=e^7x now we have to use ln because e is the base of a natural log.

OpenStudy (cutiecomittee123):

I get confused when using the laws of powers.

OpenStudy (cutiecomittee123):

especially with the "e^7x"

OpenStudy (campbell_st):

you need to recognize \[\ln(e^{7x}) = 7x\] when the base of the log and the base of the exponent are the same it becomes 1 to explain using log laws its \[e^7x = \ln(e^{7x}) = 7x \times \ln(e) = 7x \times 1\]

OpenStudy (cutiecomittee123):

The options are x=ln(3)-7 x=l/e x=ln(3)/7 x=ln(150)/7(ln(50))

OpenStudy (cutiecomittee123):

@campbell_st

OpenStudy (cutiecomittee123):

so then would it be x=ln(3)/7

OpenStudy (campbell_st):

its would be

OpenStudy (cutiecomittee123):

okay cool.